A tortoise moves at 0.5 m/s and hare moves at 6m/s if they decided to have a race over a straight 500m course how far be lined the hare will the tortoise be when the hare crosses the finish line?


1
Expert's answer
2022-03-14T15:19:20-0400

Time = "\frac{distance}{speed}"

Time for hare to finish race = "\frac{500}{6}=83.33s"

Distance = speed * time

Distance covered by tortoise as hare crosses finishing line = 0.5 * 83.33 = 41.665m

Distance between hare and tortoise = 500 - 41.665 = 458.335m

The tortoise is 458.335m behind hare.